# Model for the Thermodynamics of Iron at High Pressures Near Melting

### v2

@@ Abstract
The Fe pressure-temperature phase diagram and its melting line have a wide range of applications, including providing constraints for iron-core planetary models. We propose an equation of state (EOS) model based on the interstitial theory of simple condensed matter (ITCM) suggested by A.V. Granato [1]. When applied to Fe, this model allows for the extrapolation of measured melting lines to Earth’s inner core boundary (ICB) conditions. The ITCM describes the solid-liquid phase transition in metals as resulting from a strong structural perturbation due to a high concentration of interstitial-like defects. The nonlinearity of their self-interaction leads to the stabilization of this interstitial-rich phase. The original model is expanded to describe melting over a wide range of pressures and temperatures rather than focusing on a specific isobaric transition. The model is fitted to measured melting data, and then extrapolated to cover ICB conditions. The suggested application of ITCM naturally leads to the suggestion of two distinct liquid phases, namely HCP- and FCC-based liquids. We show that these phases and the transition between them are consistent with previous thermodynamic measurements. The additional liquid phase offers a new interpretation of the previously suggested near-melting high-pressure phase [2] and may also explain the inner core nucleation paradox [3].
@@ III.2 Liquid Phases
To maintain thermodynamic consistency, Richard’s rule [58] was incorporated, governing the entropy change during melting. Additionally, bulk and shear moduli were independently fitted to a second-order pressure and first-order temperature expressions based on ab initio calculations [59, 60, 61], allowing for extrapolation beyond the regime where measurements of the solid phases are available.
@@ IV Results and discussion
Additional Shock-Ramp melting line experiments were conducted at the National Ignition Facility (NIF) by Kraus et al. [64]. They determined the phase at each measurement point from the momentum transfer spectra. Subsequent ab initio work [28] observed that a measurement interpreted as a liquid showed deviations from classical liquid theory at both low and high frequencies, leading to an attribution of the data to a BCC phase. Instead, we propose that these deviations could be explained by resonant and local vibrational modes, which are expected within ITCM.
In addition, seismic velocity measurements and simulations have shown that Earth’s inner core displays strong shear softening and an ultrahigh Poisson’s ratio. In simulations of Fe HCP solid close to melting, this softening was attributed to the formation of increased correlation in the displacement of atoms, manifested in chains of collective motion.[65] Such increased correlation in displacement chains was previously reported in atomistic simulations of Cu’s highly perturbed crystalline phase as melting was approached.[40] Thus, this observed softening may be related to the suggested intermediate $L_{HCP}$. The formation of such an intermediate liquid phase is expected to reduce the nucleation barrier due to its effect on the solid-liquid interface energy[41, 42]. This, in turn, can explain the gap between the observed structure and the estimated nucleation time of the solid phase in Earth’s inner core. [3]. Furthermore, the $L_{HCP}$ phase is a natural candidate as an intermediate phase for a two-step process, which was previously suggested to explain this phenomenon [66].
@@ Appendix B Liquid Fitting Parameters
$L_{FCC}$ 55 1.1 0.035 0.041 0.1 4 0.28
$L_{HCP}$ 55 0.97 0.0005 0.37 0.14 10 0.07

### v3

@@ Abstract
Iron phase diagram and specifically, its melting line at high pressures, have a wide range of applications, including the study of Earth’s inner core boundary (ICB). We propose an equation of state (EOS) model based on the interstitial theory of simple condensed matter[1] (ITCM) focused on melting at high pressures. Melting is described as resulting from a strong structural perturbation due to interstitial-like defects. The nonlinearity of their interaction leads to the stabilization of interstitial-rich phase. The model is expanded to describe melting over a wide range of pressures. The model is fitted to measured melting data and extrapolated to cover ICB conditions. The ITCM naturally yields two distinct liquid phases: HCP and FCC-based liquids. These phases and the transition between them are consistent with previous measurements. The additional liquid phase offers a new interpretation of the previously suggested near-melting high-pressure phase [2] and the inner core nucleation paradox [3].
@@ III.2 Liquid Phases
To maintain thermodynamic consistency, Richard’s rule [58] was incorporated, governing the entropy change during melting. Additionally, bulk and shear moduli were independently fitted to a second-order pressure and first-order temperature expressions based on ab initio calculations [59, 60, 61], allowing for extrapolation beyond the regime where measurements of the solid phases are available. We note that the validity of this EOS depends on both the extrapolation of the solid reference states’ EOS and the model for the solid-liquid transition. Utilizing well-established analytical forms derived from physical consideration justifies the extrapolation of the free energies into the liquid regime. The transition into the liquid state addresses forming a perturbed state per solid phase. However, a not well-defined upper limit on the validity of the model stems from the fact that the properties of the reference solid state are extrapolated to an untested regime and that we do not address the formation of new solid phases, such as a high BCC phase [30, 28]
@@ IV Results and discussion
Additional evaluations of melting at high pressures was done using Shock-Ramp experiments at the National Ignition Facility (NIF) by Kraus et al. [64]. In these experiments, the phase at each measurement point was identified from the momentum transfer spectra, and deviations from classical liquid spectra were observed in a region near melting. Using ab initio simulations of Iron it was suggested that the observed signal can be explained in terms of partial formation of a BCC phase at high temperature and pressure [28]. ITCM leads to an alternate hypothesizes, where such deviations may result from increase in resonant and local vibrational modes which may be observed near melting. However this needs to be further explored using computational and experimental study.
In addition, seismic velocity measurements and simulations have shown that Earth’s inner core displays strong shear softening and an ultrahigh Poisson’s ratio. In simulations of Fe HCP solid close to melting, this softening was attributed to the formation of increased correlation in the displacement of atoms, manifested in chains of collective motion.[65] Such increased correlation leading to the formation of displacement chains was previously reported in atomistic simulations of Copper highly perturbed crystalline phase as melting was approached.[40] Thus, this observed softening may be related to the suggested intermediate $L_{HCP}$. The formation of such an intermediate liquid phase is expected to reduce the nucleation barrier due to its effect on the solid-liquid interface energy[41, 42]. This, in turn, can explain the gap between the observed structure and the estimated nucleation time of the solid phase in Earth’s inner core. [3]. Furthermore, the $L_{HCP}$ phase is a natural candidate as an intermediate phase for a two-step process, which was previously suggested to explain this phenomenon [66].
@@ V Conclusions
VI Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this manuscript.
@@ Appendix B Liquid Fitting Parameters
$L_{FCC}$ $55\pm 5$ $1.10\pm 0.02$ $035\pm 0.006 0.0$$041\pm 0.001$ $0.101\pm 0.007$ $4.00\pm 0.02$ $0.28\pm 0.015$
$L_{HCP}$ $54.9\pm 3.5$ $97\pm 0.01 0.0$$0005\pm 0.00002$ $0.371\pm 0.005$ $0.14\pm 0.01$ $10.00\pm 0.01$ $0.07\pm 0.01$
The errors presented here are estimates based on fitting the parameters of the free energies regarding the solid-liquid transition. These error estimates are partial as they do not employ a Hessian covering the complete phase diagram; developing such an estimate is beyond the scope of the current work. The error estimates were calculated by varying one parameter at a time and identifying the shift required to increase $\chi^{2}$ by one unit as illustrated in Fig. 9. This procedure provides only a sensitivity estimate as the error distribution is not Gaussian, and parameter correlations are neglected. Moreover, the evaluation is confined to the vicinity of the identified minimum. Consequently, the quoted $2\sigma$ ranges should be considered indicative rather than strict confidence intervals.
Figure 9: Illustration of the uncertainty evaluation procedure. Shown are $\chi^{2}$ variations with respect to four $L_{FCC}$ parameters ($\alpha_{1},\alpha_{2},a,b$); the same analysis was applied to all liquid-phase parameters. Uncertainties were estimated by identifying the parameter shift corresponding to an increase of $\chi^{2}$ by one unit, referred to here as the $2\sigma$.
Appendix C Data Availability
All data used in this manuscript are provided in [67].